# Saturability of the Quantum Cramér-Rao Bound in Multiparameter Quantum Estimation at the Single-Copy Level

### v1

@@ 4 Main results and discussion
The remaining steps of the proof are as follows. It will first be shown that condition (11) of Theorem 2 implies that $L_{\theta_{m},+0}L_{\theta_{l},+0}^{{\dagger}}-L_{\theta_{l},+0}L_{\theta_{m},+0}^{{\dagger}}=0$ for all $l,m=1,\ldots,p$, thus Condition 2 of the theorem is necessary. By (13) this then implies that Condition 1 of the theorem is also necessary. Note that by Lemma 7, the additional conditions in the theorem imply Condition 2. It will then be shown that Condition 1 and the additional condition stated in theorem are sufficient by showing that they guarantee the existence of a projective POVM such that condition (10) of Theorem 2 holds.
@@
Let $E_{k,00}$ have the spectral decomposition $E_{k,00}=V_{k}D_{k}V^{{\dagger}}_{k}$, where $D_{k}$ is diagonal and $V_{k}$ is unitary. Now, let $D^{+}_{k}$ be a diagonal matrix such that $D_{k}^{+}D_{k}=D_{k}^{1/2}$, such a matrix is trivial to construct. Define the matrix $H_{k}=V_{k}D^{+}_{k}V^{{\dagger}}_{k}$. Then we have that $H_{k}E_{k,00}=E_{k,00}^{1/2}$, where $E_{k,00}^{1/2}\geq 0$ is the unique Hermitian square root of $E_{k,00}$. Multiplying both sides of (14) on the left by $H_{k}$ gives
$E_{k,00}^{1/2}L_{\theta_{l},+0}^{{\dagger}}=c_{lm}^{k}E_{k,00}^{1/2}L_{\theta_{m},+0}^{{\dagger}}\quad\forall l,m=1,\ldots,p.$
From this it follows that
$L_{\theta_{m},+0}(E_{k,00}^{1/2})^{{\dagger}}E_{k,00}^{1/2}L^{{\dagger}}_{\theta_{l},+0}=L_{\theta_{l},+0}(E_{k,00}^{1/2})^{{\dagger}}E_{k,00}^{1/2}L^{{\dagger}}_{\theta_{m},+0}\quad\forall l,m=1,\ldots,p.$
Since $\sum_{k=1}^{M}(E_{k,00}^{1/2})^{{\dagger}}E_{k,00}^{1/2}=\sum_{k=1}^{M}E_{k,00}=I_{r_{0}}$, by summing the equation above over $k$ on both sides we obtain:
$L_{\theta_{m},+0}L^{{\dagger}}_{\theta_{l},+0}=L_{\theta_{l},+0}L^{{\dagger}}_{\theta_{m},+0}\quad\forall l,m=1,\ldots,p.$
Therefore Condition 2 of the theorem is necessary.
We now show that if in addition the conditions relating the corresponding columns of $L_{\theta_{l},+0}$ and $L_{\theta_{m},+0}$ for $l,m=1,\ldots,p$ as stated in the theorem are satisfied, then there exists a projective measurement with null operators that satisfy the condition (14). To this end, define $E_{k,00}$ to be a projection operator that is 0 everywhere except for a 1 at the $k$-th row and $k$-th column for $k=1,\ldots,r_{0}$. By the given conditions there is a real constant $c^{k}_{lm}$ such that the $k$-th row of $L_{\theta_{l},+0}^{{\dagger}}-c_{lm}^{k}L_{\theta_{m},+0}^{{\dagger}}$ vanishes. It follows from the construction of $E_{k,00}$ that (14) holds. Also, by construction $\sum_{k=1}^{r_{0}}E_{k,00}=I_{r_{0}}$.
Now, we turn to Condition 1. It follows from the above that this condition is necessary. It will now be shown that under Condition 1, a regular POVM operator corresponding to a projective measurement can be constructed that satisfies (10). By Lemma 4, we seek a regular POVM operator $E_{k}$ of the form,
@@
Since $[L_{\theta_{l},++},L_{\theta_{m},++}]=0$ by Condition 1, there is common set of projection operators such that $L_{\theta_{l},++}$ has the spectral decomposition
$L_{\theta_{l},++}=\sum_{k=1}^{\chi_{\theta}}\lambda_{lk}\Pi_{\theta,k},\;l=1,\ldots,p.$
where $\chi_{\theta}\leq r_{+}$ and $\Pi_{\theta,k}$ are mutually commuting projection operators. $[\Pi_{\theta,k},\Pi_{\theta,j}]=0$ and $\Pi_{\theta,j}^{2}=\Pi_{\theta,j}$ for all $j,k=1,\ldots,\chi_{\theta}$ such that $\sum_{k=1}^{\chi_{\theta}}\Pi_{\theta,k}=I_{r_{+}}$. By setting $E_{k,++}=\Pi_{\theta,k}$ for $k=1,\ldots,\chi_{\theta}$, we have that (15) is satisfied with $c^{k}_{j}=\lambda_{jk}$, $E_{k,++}\geq 0$ and $E_{k,++}\neq 0$ as required. By construction, the operator $E_{k}$ satisfies (10) and is a projection operator. Moreover, by construction, $\sum_{k=1}^{\chi_{\theta}}E_{k,++}=I_{r_{+}}$. Finally, let $E_{k}$ for $k=\chi_{\theta}+1,\ldots,\chi_{\theta}+r_{0}$ be the null POVM operators constructed earlier in the proof. It follows that $\sum_{k=1}^{M}E_{k}=I_{n_{s}}$ for $M=\chi_{\theta}+r_{0}$ as required for a POVM. This completes the proof.
The necessary conditions of the theorem are quite stringent. The first requires that $L_{\theta_{l},++}$ and $L_{\theta_{m},++}$ commute on the support subspace $\mathcal{H}_{+,\theta}$ for all $l,m=1,\ldots,p$. Besides this, the second condition must also be satisfied. The role of this condition is to reduce partial commutativity to the first condition. The following remark is also pertinent:
Remark 8
When $r_{0}=n_{s}-1$ ($r_{+}=1$), Conditions 1 and 2 in Theorem 2 are necessary and sufficient. Condition 1 holds trivially and, by (13), the necessary and sufficient average commutativity condition for pure states [15] holds if and only if Condition 2 also holds. On the other hand, when $r_{0}=1$ ($r_{+}=n_{s}-1$) then Condition 2 is not sufficient. In particular, when one of $L_{\theta_{l},+0}$ or $L_{\theta_{m},+0}$ vanishes but the other does not, Condition 2 is fulfilled but it is readily inspected that (14) only has the solution $E_{k,00}=0$.
@@
Using (6) it follows from the above identities (since $\rho_{\theta,++}>0$) that $L_{\theta_{1},+0}=(c_{2}/c_{1})L_{\theta_{2},+0}\neq 0$ and $L_{\theta_{2},++}=0$, respectively. This is enough to verify Conditions 1 and 2 and the sufficient conditions of Theorem 6 for all $\theta\in\Theta$. Therefore, the QCRB can be saturated for this quantum state by the projective measurement specified in the theorem.
@@ 5 Conclusion
This paper has established, for finite-dimensional density operators, new necessary conditions for saturability of the QCRB in single-copy multiparameter estimation that imply partial commutativity and also become sufficient with the addition of another condition. A measurement that saturates the QCRB when the sufficient conditions are satisfied is also explicitly characterized, which turns out can always be chosen to be projective. As such the results make a significant advance towards understanding conditions for saturating the QCRB in the single-copy setting and the open problem of the existence of necessary and sufficient conditions based on properties of the SLDs [9].

### v3

@@ 4 Main results and discussion
The remaining steps of the proof are as follows. It will first be shown that Conditions 1 and 2 of the theorem are necessary and sufficient. By (12) this then implies that Condition 3 is also necessary. It will then be shown that Conditions 1 and 4 (Condition 3 is then implied, Remark 7) are sufficient for the existence of a projective POVM such that (10) and (11) hold, by explicitly constructing the POVM.
First recall from §2 that $\rho_{\theta,++}$ is diagonal in the basis $\mathcal{B}_{+,\theta}$ and from (1) is given in this basis as $\rho_{\theta,++}=\mathrm{diag}(q_{1,\theta},\ldots,q_{r_{+},\theta})$, where $\mathrm{diag}(a_{1},\ldots,a_{n})$ denotes a diagonal matrix with $a_{j}$ in row and column $j$. Introduce the shorthand notation $\partial_{l}x=\partial x/\partial\theta_{l}$ and $|\partial_{l}\psi_{k,\theta}\rangle=\partial_{l}|\psi_{k,\theta}\rangle$. Differentiating,
$\partial_{l}\rho_{\theta,++}=\mathrm{diag}\left(\partial_{l}q_{1,\theta},\ldots,\partial_{l}q_{r_{+},\theta}\right).$
Let $V_{\theta}=[\begin{array}[]{ccc}|\psi_{1,\theta}\rangle&\ldots&|\psi_{r_{+},\theta}\rangle\end{array}]$. The right hand side (RHS) of the first equation in (6) can be expressed as,
$\left.P_{+,\theta}\partial_{l}\rho_{\theta}P_{+,\theta}\right|_{\mathcal{H}_{+,\theta}}$ $=\left.V_{\theta}V_{\theta}^{{\dagger}}\partial_{l}(V_{\theta}\rho_{\theta,++}V_{\theta}^{{\dagger}})V_{\theta}V_{\theta}^{{\dagger}}\right|_{\mathcal{H}_{+,\theta}}$
$=\left.V_{\theta}(V_{\theta}^{{\dagger}}\partial_{l}V_{\theta}\rho_{\theta,++}+\partial_{l}\rho_{\theta,++}+\rho_{\theta,++}\partial_{l}V_{\theta}^{\dagger}V_{\theta})V_{\theta}^{{\dagger}}\right|_{\mathcal{H}_{+,\theta}}$
$=V_{\theta}^{{\dagger}}\partial_{l}V_{\theta}\rho_{\theta,++}+\partial_{l}\rho_{\theta,++}+\rho_{\theta,++}\partial_{l}V_{\theta}^{{\dagger}}V_{\theta},$ (13)
where the last line gives the matrix representation of $\left.P_{+,\theta}\partial_{l}\rho_{\theta}P_{+,\theta}\right|_{\mathcal{H}_{+,\theta}}$ with respect to the basis $\mathcal{B}_{+,\theta}$.
For regular POVM operators $\{E_{k}\}_{k=1,\ldots,K}$ let $p_{k,\theta}=\mathrm{tr}(E_{k}\rho_{\theta})=\mathrm{tr}(E_{k,++}\rho_{\theta,++})$. Then $\{E_{k,++}\}_{k=1,\ldots,K}$ may be viewed as a POVM on $\mathcal{H}_{+,\theta}$ since they satisfy $\sum_{k=1}^{K}E_{k,++}=I_{r_{+}}$ and the probability $p_{k,\theta}$ only depends on $E_{k,++}$. Since $\rho_{\theta,++}>0$, a well-known result is that for a density operator $\rho^{\prime}_{\theta}>0$ there is a POVM that saturates the QCRB if and only if the SLDs $\{L^{\prime}_{\theta_{j}}\}_{j=1,\ldots,p}$ satisfying $(L^{\prime}_{\theta_{j}}\rho^{\prime}_{\theta}+\rho^{\prime}_{\theta}L^{\prime}_{\theta_{j}})/2=\partial_{j}\rho^{\prime}_{\theta}$ for $j=1,\ldots,p$ are mutually commuting, $[L^{\prime}_{\theta_{l}},L^{\prime}_{\theta_{m}}]=0$ for $l,m=1,\ldots,p$ [12, ref. [5] and Appendix B.1]. The RHS of the first equation of (6) does not give $\partial_{j}\rho_{\theta,++}$ but rather the RHS of (13). However, the representation basis in $\mathcal{H}_{+,\theta}$ can be freely changed from $\mathcal{B}_{+,\theta}$ to a new basis $\mathcal{C}_{+,\theta}=\{|\widetilde{\psi}_{1,\theta}\rangle,\ldots,|\widetilde{\psi}_{r_{+},\theta}\rangle\}$ defined by $[\begin{array}[]{ccc}|\widetilde{\psi}_{1,\theta}\rangle&\ldots&|\widetilde{\psi}_{r_{+},\theta}\rangle\end{array}]=V_{\theta}U^{{\dagger}}_{\theta}$, for some unitary $U_{\theta}\in\mathbb{C}^{r_{+}\times r_{+}}$. Let $\varrho_{\theta,++}=U_{\theta}\rho_{\theta,++}U_{\theta}^{{\dagger}}$ and $\widetilde{L}_{\theta_{l},++}=U_{\theta}L_{\theta_{l},++}U_{\theta}^{{\dagger}}$, using (13) we obtain from (6),
$(\widetilde{L}_{\theta_{l},++}\varrho_{\theta,++}+\varrho_{\theta,++}\widetilde{L}_{\theta_{l},++})/2$ $=U_{\theta}V_{\theta}^{{\dagger}}\partial_{l}V_{\theta}\rho_{\theta,++}U_{\theta}^{{\dagger}}+U_{\theta}\partial_{l}\rho_{\theta,++}U_{\theta}^{{\dagger}}+U_{\theta}\rho_{\theta,++}\partial_{l}V_{\theta}^{{\dagger}}V_{\theta}U_{\theta}^{{\dagger}}.$
Finally, to write the RHS of the previous equality as
$\partial_{l}U_{\theta}\rho_{\theta,++}U_{\theta}^{{\dagger}}+U_{\theta}\partial_{l}\rho_{\theta,++}U_{\theta}^{{\dagger}}+U_{\theta}\rho_{,\theta,++}\partial_{l}U_{\theta}^{{\dagger}}=\partial_{l}\varrho_{\theta,++},$
we must have that
$U_{\theta}^{{\dagger}}(\partial_{l}U_{\theta}-U_{\theta}V_{\theta}^{{\dagger}}\partial_{l}V_{\theta})\rho_{\theta,++}+\rho_{\theta,++}(\partial_{l}U_{\theta}-U_{\theta}V_{\theta}^{{\dagger}}\partial_{l}V_{\theta})^{{\dagger}}U_{\theta}$ $=0.$ (14)
Therefore, the commutativity $[L_{\theta_{l},++},L_{\theta_{m},++}]=U_{\theta}^{{\dagger}}[\widetilde{L}_{\theta_{l},++},\widetilde{L}_{\theta_{m},++}]U_{\theta}=0$ for $l,m=1,\ldots,p$ and the existence of a unitary $U_{\theta}$ satisfying (14) for $l=1,\ldots,p$ are necessary and sufficient. From (12) it follows that $L_{\theta_{l},+0}L_{\theta_{m},+0}^{{\dagger}}-L_{\theta_{m},+0}L_{\theta_{l},+0}^{{\dagger}}=0$ for all $l,m=1,\ldots,p$, thus Condition 3 of the theorem is also necessary.
@@
Let $W\in\mathbb{C}^{r_{0}\times r_{0}}$ be a unitary matrix such that the conditions on the corresponding columns of $L_{\theta_{l},+0}W$ and $L_{\theta_{m},+0}W$ for $l,m=1,\ldots,p$ as stated in Condition 4 of the theorem are satisfied. Consider $E_{k,00}=WD_{k,00}W^{{\dagger}}$ for some diagonal $D_{k,00}$ with $D_{k,00}\geq 0$. Then (15) is equivalent to
$D_{k,00}((L_{\theta_{l},+0}W)^{{\dagger}}-c^{k}_{lm}(L_{\theta_{l},+0}W)^{{\dagger}})=0,$ (16)
for all $l,m=1,\ldots,p$. By setting $D_{k,00}$ as specified in the theorem statement and $c_{lm}^{k}=\lambda_{lmk}$, with $\lambda_{lmk}$ as defined in Condition 4, $\{E_{k,00}\}_{k=1,\ldots,r_{0}}$ forms a set of projective and mutually commuting operators with $\sum_{k=1}^{r_{0}}E_{k,00}=I_{r_{0}}$. Hence the conditions are sufficient for (15).
Now, we consider Condition 1. It will be shown that under Condition 1, a regular POVM operator corresponding to a projective measurement can be constructed that satisfies (10). By Lemma 4, consider a regular POVM operator $E_{k}$,
@@
Since $[L_{\theta_{l},++},L_{\theta_{m},++}]=0$ by Condition 1, there is a common set of projection operators such that $L_{\theta_{l},++}$ has the spectral decomposition $L_{\theta_{l},++}=\sum_{k=1}^{\chi_{\theta}}\lambda_{lk}\Pi_{\theta,k},\;l=1,\ldots,p$. where $\chi_{\theta}\leq r_{+}$ and $\Pi_{\theta,k}$ are mutually commuting projection operators. $[\Pi_{\theta,k},\Pi_{\theta,j}]=0$ and $\Pi_{\theta,j}^{2}=\Pi_{\theta,j}$ for all $j,k=1,\ldots,\chi_{\theta}$ such that $\sum_{k=1}^{\chi_{\theta}}\Pi_{\theta,k}=I_{r_{+}}$. By setting $E_{k,++}=\Pi_{\theta,k}$ for $k=1,\ldots,\chi_{\theta}$, we have that (17) is satisfied with $c^{k}_{j}=\lambda_{jk}$, $E_{k,++}\geq 0$ and $E_{k,++}\neq 0$ as required. By construction, the operator $E_{k}$ satisfies (10) and is a projection operator. Moreover, by construction, $\sum_{k=1}^{\chi_{\theta}}E_{k,++}=I_{r_{+}}$. Finally, let $E_{k}$ for $k=\chi_{\theta}+1,\ldots,\chi_{\theta}+r_{0}$ be the null POVM operators constructed in the previous paragraph. It follows that $\sum_{k=1}^{M}E_{k}=I_{n_{s}}$ for $M=\chi_{\theta}+r_{0}$ as required for a POVM. This completes the proof.
The conditions of the theorem are quite stringent. Condition 1 requires that $L_{\theta_{l},++}$ and $L_{\theta_{m},++}$ commute on the support subspace $\mathcal{H}_{+,\theta}$ for all $l,m=1,\ldots,p$. Condition 2 involves a system of coupled nonlinear partial differential equations that are required to have a unitary solution $U_{\theta}$. Condition 3 is also necessary and is imposed by partial commutativity. When $r_{+}=1$, Conditions 1 and 3 in Theorem 6 are necessary and sufficient. In this case Condition 1 holds trivially and, by (12), the necessary and sufficient average commutativity condition for pure states [16] holds if and only if Condition 3 also holds. The next example illustrates a multiparameter quantum state that satisfies the conditions of the theorem.
@@
Using (6) it follows from the above identities (since $\rho_{\theta,++}>0$) that $L_{\theta_{1},+0}=(c_{2}/c_{1})L_{\theta_{2},+0}\neq 0$ and $L_{\theta_{2},++}=0$, respectively. This is enough to verify Conditions 1 and 4 (the latter with $W=1$) of Theorem 6 (hence also Condition 3 by Remark 7) for all $\theta\in\Theta$. Therefore, the QCRB can be saturated for this quantum state by the projective POVM specified in the theorem. This concludes the example.
@@ 5 Conclusion
This paper has established, for finite-dimensional density operators, necessary and sufficient conditions for saturability of the QCRB in multiparameter single-copy estimation. In addition, new necessary conditions that imply partial commutativity and that become sufficient with the addition of another condition have also been obtained. The paper explicitly characterizes the measurement that saturates the QCRB when the sufficient conditions are satisfied, which turns out can always be chosen to be projective. As such the results make a significant advance towards understanding conditions for saturating the QCRB in the single-copy setting and, in particular, resolves the open problem of necessary and sufficient conditions [10].